As an outcome of reviewing calibration results for the beam-beam counters it became clear to me that the module 20 and 23 in the right array did not match the Adc signals. (The adc spectrum did not have a sharp cutoff from the effect of the discriminator). Todays access allowed me to check the hardware and indeed the timing cable from the tubes were interchanged (the cable number 56 and 59 so you can guess someone stood upside down last time these were changed.) I talked to Hiro and this effect shouls not influence the dN/eta analysis. since the tdc were not used (for DN/deta) The cables were interchanged so from today these channels should be ok. To be able to use these also in previous runs (the problem probably has existed all through the 2001/2002 run) it is present in 4693 and in 5840 up to 6061 i modified BrRawDataInput so it will swap the raw data content. This is the only place it can be done with certainty. In later stages hits may or may not be there. So brat was updated for this today 12/13 - It is not tagged etc, since CH has announced a major update tomorrow anyhow. A new calibration were committed to the DB for run 5840. I think it is quite likely that additional calibrations has to be done for a selected set of earlier runs- not only for the gain, but also later timing offsets and slewings to gain these two channels. Morale: Do not overlook the importance of checking rawdata and results of automated procedures. ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2b30 : Thu Dec 13 2001 - 17:15:51 EST